Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day01


Crew 102 Log Book for March 26, 2011

Photos of the day






Commander's Report

Franco Carbognani Reporting

During the handover from crew 101 those are the problems found:

1) Hab

- On the outer airlock door the main window was absent. Crew 101 says it was blown off by the wind of the last few days. I have taken it, repaired  and put back in place firmly.

- The gray water system is not working and buckets must be used. In our rotation we absolutely need to have gray water for our experiments. We ask the permission to fix the system. Were exactly is the problem?

2) Suits

Pack number 5 not working

Helmets: Basically all of them are broken, in particular with the transparent part fissured or with holes (photos will follow)

3) Kitchen: The bread machine is missing the main internal tool

4) The wind has blown off the roof of the Engineering Shed. Since it was raining we have put some protection on top of the generator Kitty. We ask the permission to put the roof back in place in order to protect the engineering equipment.

5) ATV

Viking I: Does not start

Spirit: Shift gear broken


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level: Unknown

o V everstart

o V armor plate

o V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): 1 generator run


Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.20

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Notes/Comments: Propane Tank needs to be refilled


Oil:

• Kitty Oil Quantity (# of Quarts in storage): Changed/replaced by DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:


Water Status:

• Outside Potable Water Tank Level (inches from bottom): 10

• Trailer Potable Water Tank Level (inches from bottom): full

• Hab Potable Water Tank Level (inches from bottom): 8.5

• Potable Water Meter Reading: 438802

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: unchanged

• Condition of plants in Tank 3: unchanged

• Greenhab notes/ comments: can see sand, but it is still covered in a bit of water


Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

• Used (yes/no): yes

• Oil Checked (yes/no): No

• Tire Status: Not checked, good

• Odometer at end of day: 123760.0

• Notes/Comments on Hab Car:


ATV 1 (Opportunity):

• Used: no

• Oil Checked: no

• Fuel Consumed: no

• Tire Status: ok

• Comments:


ATV 2 (Spirit):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:


ATV 3 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:


ATV 4 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:

ATV Comments:


Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 70

• Actual temperature upstairs (Farenheit): 67.6

• Thermostat setting downstairs (Farenheit): 57

• Actual temperature downstairs (Farenheit): 55

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:

Webcams are not updating


General Engineering:


Journalist Report

Kavya Manyapu Reporting

On March 25th, crew 102 launched off Earth and landed on Mars on 26th March 2011. The crew landed in good health and the journey has been safe. The crew was received by the previous crew commander (crew 101) at the MDRS habitat. The whole crew was debriefed with all the operational protocols including engineering, science, safety and health, astronomy and EVA. The previous crew was then given a warm send off by Crew 102 after the hand-over. The crew received food and other logistic supplies for a 15 day survival on Analog Mars. Day 1 continued with a few engineering fixes to keep the habitat intact and warm for the rest of the mission. The crew celebrated their arrival on the Red Planet and has planned the next day operations. The crew is eagerly looking forward for a great mission, conducting various Science, Engineering and Outreach experiments during their mission. Go MDRS 102!